# NVS-SQA: Exploring Self-Supervised Quality Representation Learning for Neurally Synthesized Scenes without References

### v1

@@ IV-F Quantitative results
In addition to the numerical comparisons reported in Table X, we provide example cases in Fig. 8, where PSNR, SSIM, and LPIPS (i.e., the most widely used quality assessment methods for NSS) fail to align with actual human preferences. Each row in Fig. 8 presents a pair of neurally synthesized scenes (NSS) generated by different Neural View Synthesis (NVS) methods, along with human and NVS-SQA preferences, compared against the outcomes from PSNR, SSIM, and LPIPS.
For instance, in the top row (“Giraffe”), although PSNR and SSIM favor the right-hand result, perceptual inspection reveals substantial blurring and missing object details. By contrast, NVS-SQA and human observers correctly identify the left-hand image as exhibiting superior perceptual fidelity. A similar discrepancy is evident in the “Intr-Animals” example, where the left-hand NSS contains clear artifacts (outlined in red boxes) that standard metrics overlook, yet NVS-SQA accurately detects. In the “Orchids” scene, inconsistent color hues across views go largely unnoticed by PSNR, SSIM, or LPIPS but are recognized by our proposed method as detrimental to perceptual quality.
@@
Figure 8: Example NSS generated by various NVS methods. In all cases, the proposed NeRF-NQA aligns with human preferences without references, whereas the prevalent full-reference methods PSNR, SSIM, and LPIPS fall short. (zoom in for a clearer view)
V Conclusion

### v3

@@ IV-F Quantitative results
In addition to the numerical comparisons reported in Table X, we provide example cases in Fig. 9, where PSNR, SSIM, and LPIPS (i.e., the most widely used quality assessment methods for NSS) fail to align with actual human preferences. Each row in Fig. 9 presents a pair of neurally synthesized scenes (NSS) generated by different Neural View Synthesis (NVS) methods, along with human and NVS-SQA preferences, compared against the outcomes from PSNR, SSIM, and LPIPS. For instance, in the top row (“Giraffe”), although PSNR and SSIM favor the right-hand result, perceptual inspection reveals substantial blurring and missing object details. By contrast, NVS-SQA and human observers correctly identify the left-hand image as exhibiting superior perceptual fidelity. A similar discrepancy is evident in the “Intr-Animals” example, where the left-hand NSS contains clear artifacts (outlined in red boxes) that standard metrics overlook, yet NVS-SQA accurately detects. In the “Orchids” scene, inconsistent color hues across views go largely unnoticed by PSNR, SSIM, or LPIPS but are recognized by our proposed method as detrimental to perceptual quality.
TABLE VIII: Comparison of model complexity and inference time. The table shows the number of parameters and the average inference time per NSS for representative quality assessment methods from different categories.
Type Method Params. (M) ↓ Proc. Time per NSS (s) ↓
IQA CONTRIQUE $30.1$ $34.168$
Re-IQA $68.5$ $57.436$
VQA FAST-VQA $27.7$ 0.743
DOVER $55.5$ $0.801$
LFIQA ALAS-DADS 1.5 $1.306$
LFACon $10.5$ $0.979$
NVS-SQA (Ours) 4.7 0.692
@@
Figure 9: Example NSS generated by various NVS methods. In all cases, the proposed NeRF-NQA aligns with human preferences without references, whereas the prevalent full-reference methods PSNR, SSIM, and LPIPS fall short (zoom in for a clearer view).
IV-G Computational Cost Analysis
We present a detailed analysis of the computational efficiency and model complexity of NVS-SQA in comparison with representative no-reference IQA, VQA, and LFIQA methods. As summarized in Table VIII, NVS-SQA demonstrates the lowest inference latency, requiring only 0.692 seconds per NSS on average. In addition, it maintains a relatively compact architecture with only 4.7 million parameters, making it significantly more efficient than most existing methods in both runtime and model size.
IV-H Robustness Evaluation in Extreme Scenarios
To further examine the generalization capability of the proposed NVS-SQA model in practical conditions, we conducted a dedicated robustness evaluation under extreme scenarios that are typically challenging for novel view synthesis. These scenarios include highly sparse input views, high-noise environments, and complex lighting conditions. Such cases frequently occur in real-world applications but are often underrepresented in conventional evaluation datasets.
Evaluation setup and human label collection. We selected five challenging scenes (namely scan8, scan55, scan103, scan110, and scan114) from the DTU dataset [11], characterized by extreme conditions such as highly sparse input views, significant noise levels, and complex lighting environments. For each scene, five neural view synthesis methods were used to generate the corresponding NSS: 2DGS [70], 3D Gaussian Splatting [4], 3DLS [71], AbsGS [72], and Mip-Splatting [73]. This resulted in a total of 25 synthesized scenes across all test cases. To obtain reliable perceptual quality labels for these NSS, we conducted a user study of pairwise comparison [74, 75, 9] involving 31 participants (mean age = 25.13, SD = 1.2; 18 males, 13 females). Each participant was asked to compare ten method pairs within each scene (i.e., all unique combinations among the five methods), resulting in 50 pairwise comparisons per participant. The order of presentation and spatial placement of video sequences were randomized to eliminate bias. The collected pairwise preference data were then converted into scalar quality scores using the Bradley-Terry model [76], a statistical framework widely used for estimating latent scores from comparative data. This approach enabled the derivation of perceptual scores for each synthesized scene without requiring any reference views.
Results and Analysis. The same model weights from previous experiments were used, without any fine-tuning or adaptation. As shown in Table IX, NVS-SQA achieved a SRCC of 0.6000, a PLCC of 0.7707, and a KRCC of 0.5333. Although all evaluated methods exhibit performance degradation in extreme scenarios, the proposed method still considerably outperforms existing quality assessment methods. For example, the second-best method, FAST-VQA, achieved an SRCC of only 0.3000 and a PLCC of 0.3614. These results indicate that NVS-SQA maintains perceptual alignment with human judgments even under adverse visual conditions, effectively capturing degradations associated with sparse geometry, noise-induced artifacts, and illumination inconsistencies, without requiring any reference views or domain-specific adjustment.
TABLE IX: Robustness Evaluation under Extreme Scenarios. Performance of representative quality assessment methods on neurally synthesized scenes with sparse views, noise, and complex lighting conditions. NVS-SQA shows the highest consistency with human perception.
Type Method SRCC ↑ PLCC ↑ KRCC ↑
IQA TV 0.4732-0.4732 0.6783-0.6783 0.4300-0.4300
BRISQUE $0.0268$ $0.2121$ 0.0300-0.0300
NIQE 0.1000-0.1000 $0.0293$ 0.1333-0.1333
CLIP-IQA 0.3732-0.3732 0.5991-0.5991 0.2966-0.2966
CONTRIQUE $0.1268$ $0.0239$ $0.1034$
Re-IQA 0.1732-0.1732 0.2981-0.2981 0.1633-0.1633
VQA VIIDEO $0.1000$ $0.2550$ $0.1333$
FAST-VQA $0.3000$ $0.3614$ $0.2667$
DOVER $0.2732$ $0.3136$ $0.2966$
LFIQA ALAS-DADS $0.1217$ $0.1837$ $0.1395$
LFACon $0.2643$ $0.3215$ $0.2197$
NVS-SQA (Ours) 0.6000 0.7707 0.5333
V Limitations and Future Work
While our extensive evaluations (Section IV-H) demonstrate that NVS-SQA remains relatively stable under challenging conditions such as resolution variations and severe occlusions, it may still be limited in handling more complex motion dynamics and more diverse scene types. On the one hand, although the proposed method maintains strong correlations with human perceptual judgments, subtle degradations in scenes involving complex motion dynamics, such as rapid movements, dynamic illumination changes, or intricate texture transitions, may not always be fully captured. On the other hand, our current self-supervised training pipeline is primarily constructed using realistic 3D reconstructed scenes, which rarely contain severe synthesis artifacts found in synthetic scenes, such as large black holes, view extrapolation failures, or strong shadow distortions. This limitation may reduce the model’s sensitivity to these rare yet perceptually critical defects in synthetic scenes. In future work, we plan to enrich the training data by incorporating more diverse NVS sources and introducing controlled, manually distorted samples, particularly those exhibiting severe occlusions and shadow artifacts, to enhance the robustness and generalizability of the learned quality representations across more complex motion dynamics and diverse scene types. Another potential limitation may lie in the use of the current guidance cues, which may not yet fully exploit the complementary strengths inherent in NVS-SQA. While the proposed guidance cues (IQA, VQA, and REP) are grounded in well-established perceptual characteristics of neural scene synthesis and provide effective, complementary supervision, there may remain subtle aspects of perceptual quality that are not fully captured by these signals. Future work could investigate more flexible or data-driven supervision strategies, such as leveraging learned perceptual priors or incorporating adversarial feedback mechanisms, to further enhance the representational alignment with nuanced human perception.
VI Conclusion